OpenStudy (anonymous):

Multiply and simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. sqrt20a^7 b * sqrt8a^8 b^6

OpenStudy (anonymous):

\[ \sqrt20a^7 b * \sqrt8a^8 b^6\] square roots go all the way across

OpenStudy (anonymous):

\[\sqrt{20a^7b\times 8 a^8b^6}\] \[\sqrt{160a^{15}b^7}\] \[4a^7b^3\sqrt{10ab}\]